# Defending Multimodal Backdoored Models by Repulsive Visual Prompt Tuning

### v1

@@ 4 Experiments
Method BadNet Blended Label Consistent ISSBA WaNet TrojVQA BadCLIP
ASR CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR CA
No defense 87.49 56.34 99.22 56.08 86.14 55.97 70.19 55.69 87.58 53.05 99.72 56.62 99.83 60.63
CleanCLIP 56.55 52.77 34.33 54.22 62.97 52.74 51.85 54.06 70.03 52.37 79.69 54.90 67.53 53.42
Linear Probe 5.29 56.83 1.78 57.12 1.94 56.97 0.09 56.79 1.31 55.75 - - 99.70 59.33
RVPT 1.84 58.09 0.02 57.97 1.14 57.92 0.01 58.18 0.09 56.22 3.33 58.12 2.76 61.72
@@
ASR CA ASR CA ASR CA ASR CA ASR CA ASR CA
No defense 88.85 91.07 91.28 90.69 80.91 88.63 93.89 80.36 100 71.24 99.91 75.81
CleanCLIP 29.40 88.26 15.72 90.25 24.66 91.10 72.57 70.45 32.98 73.25 88.06 75.11
Linear Probe 1.47 92.53 0.53 93.37 1.03 91.74 9.02 75.75 3.71 70.43 3.70 74.89
RVPT 0.04 93.55 0 93.40 0.28 93.33 0.64 86.02 2.97 83.62 2.68 85.49
@@ 4.1 Experiment Settings
Models and benchmarks. Following the settings of [2], to create a more realistic backdoored CLIP for backdoor defense, we fine-tune OpenAI’s public checkpoint CLIP-400M [41] using 500K image-text pairs from the CC3M dataset [44], out of 1500 are poisoned with various types of backdoor attacks: BadNet [15], Blended [6], Label Consistent [46], ISSBA [25], WaNet [37], TrojVQA [48] and BadCLIP [28]. For BadNet, Blended, and WaNet, the target class is banana, accordion, or samoyed. For Label Consistent, ISSBA, WaNet, TrojVQA and BadCLIP, the target class is banana. Our defense mainly utilizes ImageNet [9] to defend against backdoor attacks whose target class is banana, Caltech101 [13] for accordion and OxfordPet [39] for samoyed. We compare RVPT with Linear Probe and CleanCLIP [2]. The implementation details of the aforementioned backdoor attack methods, backdoor defense methods, and benchmarks will be thoroughly illustrated in the supplemental material. Additionally, we report the zero-shot results of CleanCLIP and RVPT in the following experiment. The results of combinations of defense methods, such as CleanCLIP + Linear Probe, RVPT + Linear Probe, CleanCLIP + RVPT will also be provided in the supplemental material.
Evaluation criteria. For backdoor attacks targeting the class of banana, accordion, or samoyed, we evaluate the defense methods utilizing the test sets of ImageNet, Caltech101, and OxfordPets separately. The evaluation metrics include Clean Accuracy (CA) and Attack Success Rate (ASR). CA is the model’s accuracy on clean images, while ASR measures the proportion of poisoned images predicted as the target label. A lower ASR and decent CA stand for a successful defense.
Implementation details. The implementation of our defense method is based on Pytorch [40]. Unless otherwise specified, we will use ViT/B32 [11] as CLIP’s visual encoder. We set the non-prompted depth $D=3$, learnable context length $b=50$, and balancing factor $\alpha=2$. Moreover, we set the proxy caption of the class the same as the simple prompt engineering in the original paper of CLIP [41]: (1) “a photo of <CLS>.” for ImageNet and Caltech101 (2) “a photo of <CLS>, a type of pet.” for OxfordPets. For all datasets, we randomly sample 16 images per class while setting the batch size and the epoch number to 32 and 50. The loss is optimized with SGD with an initial learning rate of 0.002 decayed by the cosine annealing rule.
@@ 4.2 The Effectiveness of RVPT
In Table 1 and 2, we compare RVPT with CleanCLIP and Linear Probe against seven attacks on ImageNet and three attacks on Caltech101 and OxfordPets.
Firstly, we can see all attacks achieve a high ASR while maintaining decent CA with no defense, indicating all attacks have been successfully launched.
Secondly, although CleanCLIP mitigates backdoor attacks in a zero-shot setting, it deteriorates CLIP’s recognition performance on downstream tasks and requires more computing resources (it needs full parameter tuning and much more image-text pairs). Moreover, Linear Probe performs well against the first five attacks without losing clean accuracy. However, it actually does not remove the backdoor implanted in the visual encoder and performs the worst against BadCLIP, an attack specially designed for multimodal contrastive learning models. Since CLIP tuned with Linear Probe will only retain visual encoder, we opted not to apply it to defend against TrojVQA, which focused on the multimodal task of visual question answering.
Finally, RVPT successfully defends against all types of backdoor attacks, with a notable performance gain: CA increased by 1% to 2% generally. In particular, it performs better against global trigger patterns, such as Blended, ISSBA, or WaNet. We conjecture the reason behind this phenomenon lies in the fact that the global trigger pattern rarely exists in the downstream dataset, which is easier to be recognized as the class-irrelevant feature and removed by RVPT.
@@
Target class RVPT
BadNet (woTC) BadNet Blended (woTC) Blended WaNet (woTC) WaNet
banana 2.18 (58.15) 1.84 (58.09) 0.01 (57.73) 0.02 (57.97) 0.09 (55.34) 0.09 (56.22)
samoyed 0.76 (84.25) 0.64 (86.02) 0.03 (81.64) 2.97 (83.62) 1.03 (83.76) 2.68 (85.39)
accordion 2.29 (93.41) 0.04 (93.55) 0 (92.83) 0 (93.40) 0 (92.94) 0.28 (93.33)
@@
Method accordion samoyed
@@
CleanCLIP 29.40 (88.26) 15.72 (90.25) 24.66 (91.10) 72.57 (70.45) 32.98 (73.25) 88.06 (75.11)
RVPT* 5.59 (87.47) 1.02 (88.03) 2.49(89.02) 5.97 (76.90) 0.64 (76.53) 3.42 (76.79)
4.3 RVPT’s Defensive Ability Can Generalize
We evaluate RVPT’s generalization ability in defending backdoor attacks across three scenarios:
  • The target class of the attacks is not in the tuning dataset.
  • The attacks are launched for downstream datasets different from the tuning dataset.
  • The attacks are launched for the cross-domain datasets.
@@ 4.3 RVPT’s Defensive Ability Can Generalize
No defense 89.06 (49.33) 99.53 (49.04) 99.89 (53.49)
CleanCLIP 62.92 (45.58) 40.57 (47.24) 73.07 (46.20)
RVPT 2.64 (49.35) 0.02 (49.51) 3.43 (52.53)
@@
No defense 92.96 (23.31) 99.92 (24.10) 99.97 (30.80)
CleanCLIP 79.19 (19.25) 71.03 (20.79) 92.27 (20.58)
RVPT 7.67 (14.25) 0.09 (14.69) 12.84 (16.84)
@@
No defense 89.59 (61.24) 98.61 (61.97) 99.69 (65.49)
CleanCLIP 68.66 (58.79) 44.31 (60.29) 72.32 (60.29)
RVPT 3.56 (53.08) 0.08 (53.02) 6.75 (57.37)
@@
No defense 93.81 (36.43) 96.48 (36.88) 99.88 (40.19)
CleanCLIP 76.37 (33.38) 14.26 (34.33) 69.80 (34.46)
RVPT 1.92 (31.16) 0 (30.76) 1.67 (34.06)
Performance on the emerging target class.  We remove the samples belonging to the target class from the tuning dataset to evaluate RVPT’s performance on the emerging target class, which measures its open-vocabulary defense capability for unseen target classes, and the results are shown in Table 3. Moreover, we also tested BadCLIP (woTC) for the target class of banana and the ASR (CA) is 3.39 (61.89). Table 3 shows that for the backdoor attack of Blended and WaNet, RVPT performs even better in the absence of the target class, suggesting that RVPT has successfully identified and removed the distinctive pattern. For BadNet, which is harder to resist, ASR increases slightly, though still within a reasonable range. Moreover, we can see that CA maintains other than OxfordPets, which decreases by around 2%. We believe the reason is that this fine-grained dataset only contains 37 classes. The absence of 3% of the data can significantly impact accuracy. In summary, RVPT’s defensive capability effectively generalizes to attacks with emerging target classes that are not seen in the tuning dataset.
Performance on cross-dataset transfer.  To evaluate RVPT’s generalization of defensive ability across datasets, we fine-tuned RVPT on ImageNet and tested it on Caltech101 and OxfordPets. The results are shown in Table 4. For all three attacks, although ASR increases slightly, it remains at a low level. Additionally, RVPT’s accuracy decreases occasionally. We conjecture the reason is that ImageNet does not share some of the class-irrelevant features (CIFs) with Caltech101 or OxfordPets. In some cases, The CIFs in ImageNet may be the discriminative features in them. Nevertheless, the CIFs of real-world images seem quite similar, ensuring RVPT’s cross-dataset defensive ability. Moreover, VPT’s tendency to overfit the base dataset is one reason for the CA drop. Overall, RVPT’s performance on new datasets declines slightly but still defends better than CleanCLIP.
Performance on cross-domain transfer.  To evaluate the cross-domain defensive ability, RVPT was trained on ImageNet and tested on ImageNet-V2 [43], ImageNet-A [17], ImageNet-R [16], and ImageNet-S [50]. The result is shown in Table 5. These three attacks can generalize well on the four benchmarks. We observe that RVPT will experience a loss of recognition and defense performance with domain shifts, especially in ImageNet-A. We thus conjecture that the CIFs of cross-domain images vary more than real-world images. Notably, even with cross-domain images, RVPT still outperforms CleanCLIP and significantly decreases ASR.
@@ 4.4 Ablation Studies
Figure 3: ASR and CA are evaluated when the hyperparameters of shots, $\alpha$, context token length, depth, and margin are changed in RVPT.
Number of shots per class.  How many shots are needed to guarantee a successful defense? From Figure 3 (a), we surprisingly find that only 1 shot per class can make the ASR of Blended less than 0.02% and significantly decrease the ASR for other attacks. Moreover, increasing the number of shots further enhances the defense. For BadNet and BadCLIP, 16 shots are needed to bring the ASR below 5%. Nevertheless, with 32 shots, the defense effectiveness stops improving, which might be caused by the limited repelling capabilities of the visual prompts.
Balancing factor $\alpha$.  We use $\alpha$ to balance the strengths of the CE loss and FR loss. From Figure 3 (b), we observe that ASR sharply decreases while CA sharply decreases as $\alpha$ increases. This observation indicates that if the balancing factor is set too high, it will interfere with the encoding of class-relevant features (the environment is too harsh for CE loss optimization), resulting in performance degradation. However, a relatively small balancing factor, like 2, can achieve very low ASR while preserving relatively strong CA for every evaluated dataset.
Visual context length.  We show RVPT’s performance in Figure 3 (c) with changing visual context length $b$. It indicates that both ASR and CA are optimal at a context length of approximately 25 to 50 tokens. Beyond this optimal range, CA decreases due to overfitting caused by the increasing number of parameters. Moreover, with increasing length, the defensive ability is bounded.
Depth of the visual prompt.  Figure 3 (d) illustrates the performance as the depth of the features into which the visual prompts are embedded is varied. It is important to note that “depth” here refers to the number of layers counted from the model’s output feature, as the later layers are more significant for defending against backdoor attacks. This phenomenon can be observed in Figure 3 (e), which reflects the influence of depth measured from the input features. The rationale behind this observation lies in the nature of class-irrelevant features encoded in the earlier layers, which tend to be sparse and intermixed with desirable features. Consequently, repelling these class-irrelevant features at an earlier stage is difficult and may inadvertently compromise the encoding of discriminative features. Moreover, the results in Figure 3 (d) indicate the improving performance as the number of the prompted layers increases. Particularly, when the depth reaches approximately six layers, both CA and ASR attain their optimal values and will not deteriorate.
@@
A greater $m$ means a lower level of repulsion, and $m=1$ signifies no repulsion. The performance of RVPT of different $m$ is shown in Figure 3 (f). We observe that both ASR and CA are sensitive with $m$ and will increase with larger $m$. In order to achieve superior defense performance, we adopt full repulsion that optimizes the mean cosine similarity to -1.
FR loss.  To assess the backdoor robustness introduced by FR loss in deep visual prompt tuning, we compare the ASR of RVPT and VPT against BadNet attack and Blended attack in different scenarios, as shown in Figure 4. The results can also be regarded as an ablation analysis of RVPT, which indicates that while VPT alone reduces ASR to some extent, FR loss significantly enhances the defensive capability against backdoor attacks across various conditions.

### v2

@@ 4 Experiments
Method BadNet Blended ISSBA WaNet TrojVQA BadCLIP
No defense 82.69 (63.04) 98.52 (62.64) 60.01 (61.72) 87.18 (62.42) 99.75 (62.81) 99.83 (61.33)
CleanCLIP 23.79 (57.91) 0.25 (57.69) 15.62 (59.20) 11.10 (59.07) 85.64 (58.22) 89.70 (57.55)
Linear Probe 3.05 (59.64) 5.52 (59.69) 0.08 (59.69) 0.65 (59.66) - 99.70 (59.33)
RVPT 0.05 (62.76) 0.02 (62.36) 0.01 (61.92) 0.03 (62.48) 0.11 (62.63) 2.76 (61.81)
@@
No defense 91.38 (93.06) 92.69 (93.41) 63.21 (92.86) 86.83 (82.91) 99.80 (85.10) 87.97 (83.93)
CleanCLIP 36.87 (91.18) 1.14 (90.77) 9.35 (91.54) 25.72 (82.49) 4.17 (83.41) 12.61 (81.10)
Linear Probe 1.22 (93.62) 12.82 (93.41) 1.04 (93.45) 16.05 (77.74) 2.63 (77.63) 2.21 (77.71)
RVPT 0.00 (94.02) 0.00 (94.34) 0.08 (93.89) 0.30 (88.60) 0.64 (88.87) 1.59 (88.53)
@@ 4.1 Experiment Settings
Models and benchmarks. Following the settings of [3], to create a more realistic backdoored CLIP for backdoor defense, we fine-tune OpenAI’s public checkpoint CLIP-400M [45] using 500K image-text pairs from the CC3M dataset [48], out of 1500 are poisoned with various types of backdoor attacks: BadNet [17], Blended [7], ISSBA [28], WaNet [41], TrojVQA [52] and BadCLIP [31]. We utilize ImageNet [10] to evaluate the performance of defense methods against all the aforementioned attacks, with the target class of the attacks set to banana. Additionally, we use Caltech101 [14] and OxfordPet [43] (a fine-grained dataset) to evaluate defense methods against BadNet, Blended, and WaNet. These attacks target “accordion” in Caltech101 and “samoyed” in OxfordPet, respectively. We compare RVPT with Linear Probe and CleanCLIP [3]. The implementation details of the backdoor attack methods, backdoor defense methods, and benchmarks will be illustrated in the supplementary material.
Evaluation criteria. The evaluation metrics include Clean Accuracy (CA) and Attack Success Rate (ASR). CA is the model’s accuracy on clean images, while ASR measures the proportion of poisoned images predicted as the target label. A lower ASR and decent CA stand for a successful defense.
Implementation details. The implementation of our defense method is based on Pytorch [44]. Unless otherwise specified, we will use ViT/B32 [12] as CLIP’s visual encoder. We set the non-prompted depth $D=3$, learnable context length $b=50$, and balancing factor $\alpha=2$. Moreover, we set the proxy caption of the class the same as the simple prompt engineering in the original paper of CLIP [45]: (1) “a photo of <CLS>.” for ImageNet and Caltech101 (2) “a photo of <CLS>, a type of pet.” for OxfordPets. For all datasets, we randomly sample 16 images per class while setting the batch size and the epoch number to 32 and 50. The loss is optimized with SGD with an initial learning rate of 0.002 decayed by the cosine annealing rule. We conduct experiments on eight NVIDIA RTX 3090 GPUs and the computational expenses are shown in the supplementary material.
@@ 4.2 The Effectiveness of RVPT
In Table 1 and 2, we compare RVPT with CleanCLIP and Linear Probe against six attacks on ImageNet and three attacks on Caltech101 and OxfordPets. Before poisoning, CLIP’s CA on ImageNet, Caltech101, and OxfordPets is 62.01%, 91.46%, and 86.42%, respectively.
Firstly, we can see all attacks achieve a high ASR while maintaining decent CA with no defense, indicating their successful executions. The CA loss of the poisoned CLIP varies across downstream datasets because to obtain a backdoored CLIP, fine-tuning on the poisoned CC3M dataset [48] changes CLIP’s learned representations. These changes affect how CLIP performs on different datasets, accounting for varying levels of CA loss.
Table 3: Poisoned Accuracy (PA), defined as the model’s classification accuracy on poisoned images. Among these defense methods, RVPT achieves the highest PA, demonstrating its superior capability in preventing backdoored models from encoding the trigger, thereby maintaining accuracy on poisoned samples.
Model BadNet Blended BadCLIP
Clean Model 58.29 54.16 55.34
Backdoored Model 10.36 1.05 0.18
CleanCLIP 39.66 46.12 4.25
Linear Probe 28.02 14.94 5.64
RVPT 58.94 53.27 53.90
Secondly, although CleanCLIP mitigates backdoor attacks in a zero-shot setting, it deteriorates CLIP’s recognition performance on downstream tasks and requires more computing resources (it needs full parameter tuning and more image-text pairs). Moreover, Linear Probe performs well against the first four attacks. However, it does not eliminate the trigger feature and thus performs the worst against BadCLIP, an attack specially designed for multimodal contrastive learning models. Since Linear Probe inevitably excludes CLIP’s text encoder, it cannot be used to defend against TrojVQA, which targets the multimodal task of visual question answering.
Finally, RVPT effectively defends against all types of attacks while maintaining clean accuracy. It demonstrates consistently strong defensive performance across all attacks and, notably, significantly reduces the ASR of BadCLIP [31], which is hard to defend by current defense methods. Moreover, as shown in Table 3, its classification accuracy on poisoned images is also much higher than other defense methods, highlighting its enhanced capability in maintaining correct predictions under backdoor attacks. Moreover, we observed that even a non-backdoored CLIP experiences a significant drop in PA, further supporting our claim of its low visual feature resistivity to input perturbation.
4.3 RVPT’s Defensive Ability Can Generalize
We evaluate RVPT’s generalization ability in defending backdoor attacks in three scenarios:
  • The target class of the attacks is absent from the tuning dataset but appears during testing.
  • The downstream dataset that will be attacked is different from the tuning dataset.
  • The downstream dataset that will be attacked shares different domains with the tuning dataset.
@@ 4.3 RVPT’s Defensive Ability Can Generalize
Attack Type ImageNet (woTC) ImageNet Caltech101 (woTC) Caltech101 OxfordPets (woTC) OxfordPets
BadNet 0.01 (62.50) 0.05 (62.76) 0.04 (93.71) 0.00 (94.02) 0.33 (86.84) 0.30 (88.60)
Blended 0.01 (62.63) 0.02 (62.36) 0.00 (93.89) 0.00 (94.34) 0.00 (87.19) 0.64 (88.87)
WaNet 0.00 (62.46) 0.03 (62.48) 0.00 (93.67) 0.08 (93.89) 0.00 (86.17) 1.59 (88.53)
@@
Method Caltech101 OxfordPets
@@
CleanCLIP 36.87 (91.18) 1.14 (90.77) 9.35 (91.54) 25.72 (82.49) 4.17 (83.41) 12.61 (81.10)
RVPT* 2.32 (89.33) 0.61 (89.91) 1.80 (89.66) 2.86 (82.65) 0.53 (83.49) 2.49 (82.66)
@@
No defense 86.55 (55.39) 99.04 (54.83) 99.89 (53.49)
CleanCLIP 31.38 (50.95) 0.42 (50.96) 92.04 (50.60)
RVPT 0.04 (53.85) 0.02 (53.77) 3.43 (52.53)
@@
No defense 92.97 (31.47) 99.89 (31.22) 99.97 (30.80)
CleanCLIP 59.11 (25.71) 3.18 (27.10) 98.18 (25.52))
RVPT 1.64 (16.52) 0.17 (16.84) 12.84 (16.84)
@@
No defense 66.63 (67.11) 97.94 (66.06) 99.69 (65.49)
CleanCLIP 32.99 (61.81) 2.54 (60.69) 89.03 (60.93)
RVPT 0.76 (58.39) 0.09 (58.46) 6.75 (57.37)
@@
No defense 92.11 (41.73) 97.16 (41.86) 99.88 (40.19)
CleanCLIP 26.54 (34.62) 0.26 (34.82) 85.62 (34.44)
RVPT 0.02 (35.17) 0.01 (35.34) 1.67 (34.06)
Performance on the emerging target class.  We remove the samples belonging to the target class from the tuning dataset to evaluate RVPT’s performance on the emerging target class, which measures its defense capability for unseen target classes in the dataset, and the results are shown in Table 4. Moreover, we also tested BadCLIP (woTC) for the target class of banana and the ASR (CA) is 3.39 (61.89). Table 4 shows that for the backdoor attack of Blended and WaNet, RVPT performs even better in the absence of the target class, suggesting that RVPT has successfully filtered out the distinctive out-of-dataset pattern. Moreover, we can see that CA maintains other than OxfordPets, which decreases by around 2%. The CA drop is due to the dataset containing only 37 classes, so the absence of 3% of the data can considerably impact its accuracy. In summary, RVPT’s defensive capability effectively generalizes to attacks with emerging target classes that are not seen in the tuning dataset.
Performance on cross-dataset transfer.  To evaluate RVPT’s defensive ability across datasets, we fine-tuned RVPT on ImageNet and tested it on Caltech101 and OxfordPets. The results are shown in Table 5. For all three attacks, although ASR increases slightly, it remains at a very low level. Additionally, RVPT’s accuracy decreases and occasionally falls below that of CleanCLIP. We conjecture the reason is that ImageNet does not share some of the in-dataset predictive features with Caltech101 or OxfordPets. In some cases, The non-essential feature unlearned in ImageNet may be the discriminative features in them and some of their non-essential features are not removed during tuning in ImageNet. Thus, there is a slight CA loss and ASR increase in these two datasets after RVPT. Nevertheless, the in-dataset predictive features of real-world images seem quite similar, ensuring RVPT’s cross-dataset defensive ability. Moreover, VPT’s tendency to overfit the base dataset is also one reason for the CA drop. Overall, RVPT’s performance on new datasets declines slightly but still defends better than CleanCLIP.
Performance on cross-domain transfer.  To evaluate the cross-domain defensive ability, RVPT was trained on ImageNet and tested on ImageNet-V2 [47], ImageNet-A [19], ImageNet-R [18], and ImageNet-S [54]. The result is shown in Table 6. These three attacks can generalize well on the four benchmarks. We observe that RVPT will experience a loss of recognition and defense performance with domain shifts. In particular, it shows a significant CA drop on ImageNet-A, which reflects its effectiveness in focusing on in-dataset predictive features and ignoring other features, since ImageNet-A intentionally collects natural adversarial samples for ImageNet classifiers. Notably, unlike adversarial training [2], RVPT is not designed to defend against adversarial samples in ImageNet-A. The detailed comparison between RVPT and adversarial training is in the supplementary material.
@@ 4.4 Ablation Studies
Figure 3: ASR and CA are evaluated when the hyperparameters of shots, $\alpha$, context token length, depth, and margin are changed in RVPT. More ablation studies of visual encoder architecture and handcrafted text prompt can be referred to in the supplementary material.
Number of shots per class.  From Figure 3 (a), we surprisingly find that only 1 shot per class can make the ASR of Blended and BadNet attacks nearly 0 and significantly decrease the ASR of BadCLIP and increasing the number of shots further enhances the defense. For BadCLIP, 8 shots are needed to bring the ASR below 5%. Nevertheless, with 32 shots, the defense effectiveness stops improving.
Balancing factor $\alpha$.  We use $\alpha$ to balance the strengths of the CE loss and FR loss. From Figure 3 (b), we observe that both ASR and CA decrease as $\alpha$ increases. This observation indicates that if the balancing factor is set too high, it will interfere with the encoding of in-dataset predictive features (the adversarial environment is too harsh for CE loss optimization), resulting in performance degradation. However, a relatively small $\alpha$ is sufficient to achieve very low ASR while preserving strong CA for every evaluated dataset.
Visual context length.  We show RVPT’s performance in Figure 3 (c) with changing visual context length $b$. It indicates that both ASR and CA are optimal at a context length of approximately 25 to 50 tokens. Beyond this optimal range, CA decreases due to overfitting caused by the increasing number of parameters. Moreover, when the length is 5, the ASRs of all attacks drop below 5%, indicating that tuning only 0.5‱ of parameters is sufficient to defend against all attacks, demonstrating the efficiency of our method. However, with increasing length, the defensive ability is bounded.
Depth of the visual prompt.  Figure 3 (d) shows the performance as the depth of the features into which the visual prompts are embedded is varied. It is important to note that “depth” here refers to the number of layers counted from the model’s output feature. We also conduct experiments where prompts are stacked from the input feature, as shown in Figure 3 (e). The comparison demonstrates that the later layers are more significant for backdoor robustness and recognition performance. We conjecture the reason is that the non-essential features encoded in the earlier layers tend to be sparse and intermixed with desirable features and repelling these features at an earlier stage is difficult and may inadvertently compromise the encoding of discriminative features, resulting in worse performance. Therefore, unless otherwise specified, we will count the depth from the output feature. Moreover, the results in Figure 3 (d) indicate the improving performance as the number of the prompted layers increases. Particularly, when the depth reaches approximately six layers, both CA and ASR attain their optimal values.
@@
A greater $m$ means a lower level of repulsion, and $m=1$ signifies no repulsion. The performance of RVPT of different $m$ is shown in Figure 3 (f). We observe that both ASR and CA will increase with larger $m$. Therefore, for the purpose of achieving superior defense performance, we adopt full repulsion that optimizes the mean cosine similarity to -1.
FR loss.  To assess the backdoor robustness introduced by FR loss in deep visual prompt tuning, we compare the ASR of RVPT and VPT against BadNet attack and Blended attack in different scenarios, as shown in Figure 4. The result indicates that while VPT alone reduces ASR to some extent, FR loss significantly enhances the defensive capability against backdoor attacks across various conditions.